# Long-time asymptotics of (1,3)-sign solitary waves for the damped nonlinear Klein-Gordon equation

### v1

@@ 6.1. Sharp estimates for lengths and angles
Therefore, (6.19) holds for all $t$ with $0<t-T_{1}\ll 1$, which contradicts the definition (6.20). Hence $T_{1}=\infty$. In this case, $\max\{d_{12},d_{13},d_{23}\}$ is bounded and monotone increasing, so it converges to a nonzero limit; this contradicts (6.4). Therefore, for $t\gg 1$
$\max{(d_{12},d_{13},d_{23})}\lesssim\frac{1}{L^{4}}.$ (6.21)
By (6.6), (6.17) and (6.21), we obtain (6.5), and we complete the proof.
@@ 6.2. Long-time dynamics of $(1,3)$-sign solitary waves
By the previous subsection, the estimates for $\rho_{1},\rho_{2},\rho_{3}$ have been improved. Using these, we first derive a more precise system of ODEs for $z_{0},z_{1},z_{2},z_{3}$.
@@
$\begin{aligned} \dot{z}_{0}&=-\frac{\mathcal{F}(L)}{L}(Z_{1}+Z_{2}+Z_{3})+O\left(\frac{1}{t(\log{t})^{4}}\right),\\ \dot{z}_{1}&=\frac{\mathcal{F}(L)}{L}Z_{1}+O\left(\frac{1}{t(\log{t})^{4}}\right),\\ \dot{z}_{2}&=\frac{\mathcal{F}(L)}{L}Z_{2}+O\left(\frac{1}{t(\log{t})^{4}}\right),\\ \dot{z}_{3}&=\frac{\mathcal{F}(L)}{L}Z_{3}+O\left(\frac{1}{t(\log{t})^{4}}\right).\end{aligned}$ (6.22)
@@
$\left|\mathcal{F}(\rho_{k})-\mathcal{F}(L)\right|\lesssim\frac{1}{t(\log{t})^{4}},\ \left|\frac{1}{\rho_{k}}-\frac{1}{L}\right|\lesssim\frac{1}{(\log{t})^{6}}.$
This proves Lemma 6.3. ∎
@@
Let $\vec{u}$ be a (1,3)-sign solitary waves. Then, there exists $z_{g}\in\mathbb{R}^{d}$ such that we have
$\left|\frac{z_{0}(t)+z_{1}(t)+z_{2}(t)+z_{3}(t)}{4}-z_{g}\right|\lesssim\frac{1}{(\log{t})^{3}}$ (6.23)
@@
By Lemma 6.3, we have
$\left|\frac{\dot{z}_{0}+\dot{z}_{1}+\dot{z}_{2}+\dot{z}_{3}}{4}\right|\lesssim\frac{1}{t(\log{t})^{4}}.$
Since the right-hand side is integrable, there exists $z_{g}\in\mathbb{R}^{d}$ such that $\frac{z_{0}+z_{1}+z_{2}+z_{3}}{4}$ converges to $z_{g}$ as $t\to\infty$. In particular, we have
$\left|\frac{z_{0}(t)+z_{1}(t)+z_{2}(t)+z_{3}(t)}{4}-z_{g}\right|$ $\lesssim\int_{t}^{\infty}\frac{ds}{s(\log{s})^{4}}$
$\lesssim\frac{1}{(\log{t})^{3}}.$
@@
Next, we prove the convergence of $z_{0}$. The idea, as in [21, Lemma 4.8], is to argue that if $Z_{1}+Z_{2}+Z_{3}$ did not converge, then its magnitude would grow and become much larger than $\rho_{12},\rho_{13},\rho_{23}$, leading to a contradiction.
@@
Let $\vec{u}$ be a (1,3)-sign solitary waves. Then, there exists $z_{\infty}\in\mathbb{R}^{d}$ such that
$|z_{0}(t)-z_{\infty}|\lesssim\frac{1}{(\log{t})^{3}}.$ (6.24)
@@
$\begin{aligned} \dot{Z}_{1}&=\frac{\mathcal{F}(L)}{L}(2Z_{1}+Z_{2}+Z_{3})+O\left(\frac{1}{t(\log{t})^{4}}\right),\\ \dot{Z}_{2}&=\frac{\mathcal{F}(L)}{L}(Z_{1}+2Z_{2}+Z_{3})+O\left(\frac{1}{t(\log{t})^{4}}\right),\\ \dot{Z}_{3}&=\frac{\mathcal{F}(L)}{L}(Z_{1}+Z_{2}+2Z_{3})+O\left(\frac{1}{t(\log{t})^{4}}\right).\end{aligned}$ (6.25)
From (6.25), the quantities $\rho_{12},\rho_{13},\rho_{23}$ also satisfy
$\dot{\rho}_{12}$ $=\frac{\mathcal{F}(L)}{L}\rho_{12}+O\left(\frac{1}{t(\log{t})^{4}}\right),$
$\dot{\rho}_{13}$ $=\frac{\mathcal{F}(L)}{L}\rho_{13}+O\left(\frac{1}{t(\log{t})^{4}}\right),$
$\dot{\rho}_{23}$ $=\frac{\mathcal{F}(L)}{L}\rho_{23}+O\left(\frac{1}{t(\log{t})^{4}}\right).$
In particular, when we define
$\mathcal{R}=\rho_{12}+\rho_{13}+\rho_{23},$ (6.26)
we have
$\dot{\mathcal{R}}=\frac{\mathcal{F}(L)}{L}\mathcal{R}+O\left(\frac{1}{t(\log{t})^{4}}\right).$ (6.27)
We define $W=Z_{1}+Z_{2}+Z_{3}$ and $\mathcal{W}=|W|$. Then, $W$ satisfies
$\dot{W}=\frac{4\mathcal{F}(L)}{L}W+O\left(\frac{1}{t(\log{t})^{4}}\right).$
and therefore there exists $C_{1}>1$ such that for $t\geq 2$
$\left|\dot{\mathcal{W}}-\frac{4\mathcal{F}(L)}{L}\mathcal{W}\right|\leq\frac{C_{1}\mathcal{F}(L)}{L^{4}}.$ (6.28)
@@
$\mathcal{W}(t^{\prime})>\frac{C_{1}}{L^{3}(t^{\prime})}.$ (6.29)
@@
Furthermore, we assume that $T_{1}<\infty$. Then we have $\mathcal{W}(T_{1})=\frac{C_{1}}{L^{3}(T_{1})}$. Furthermore, by (6.28), we have
$\dot{\mathcal{W}}(T_{1})\geq\frac{4\mathcal{F}(L(T_{1}))}{L(T_{1})}\mathcal{W}(T_{1})-\frac{C_{1}\mathcal{F}(L(T_{1}))}{L^{4}(T_{1})}>0.$
Moreover, since $\frac{C_{1}}{L^{3}}$ is monotone decreasing, the estimate (6.30) also holds for $0<t-T_{1}\ll 1$, which contradicts the maximality of $T_{1}$. Hence $T_{1}=\infty$. Since $\mathcal{W}$ is monotone increasing for $t>t^{\prime}$, if $\mathcal{W}(t)$ did not tend to $\infty$ as $t\to\infty$, then $\mathcal{W}$ would converge. In that case, for $1\ll T_{2}\ll T_{3}$ we would have
$1\gtrsim\mathcal{W}(T_{3})-\mathcal{W}(T_{2})\gtrsim\int_{T_{2}}^{T_{3}}\frac{dt}{t\log t},$
@@
$=-\frac{3\mathcal{F}(L)}{L}\frac{\mathcal{R}}{\mathcal{W}}+O\left(\frac{\mathcal{W}+\mathcal{R}}{\mathcal{W}^{2}t(\log{t})^{4}}\right)$
$=-\frac{3\mathcal{F}(L)}{L}\mathcal{X}+O\left(\frac{1+\mathcal{X}}{t(\log{t})^{4}}\right),$
@@
which contradicts to (6.33). Thus, we obtain
$\mathcal{W}=|z_{1}+z_{2}+z_{3}-3z_{0}|\lesssim\frac{1}{L^{3}}\sim\frac{1}{(\log{t})^{3}}.$
@@
Here, we prove Theorem 1.1. First, we show the following lemma.
@@
Let $\vec{u}$ be a (1,3)-sign solitary waves. Then, there exist $z_{\infty}\in\mathbb{R}^{d}$ and equilateral-triangular $\omega_{1},\omega_{2},\omega_{3}\in S^{d-1}$ such that for $k=1,2,3$
@@
$\dot{Z}_{k}$ $=\mathcal{F}(\rho_{k})u_{k}+O\left(\frac{1}{t(\log{t})^{4}}\right),$
$\dot{\rho}_{k}$ $=\mathcal{F}(\rho_{k})+O\left(\frac{1}{t(\log{t})^{4}}\right),$
and we have
@@
$=O\left(\frac{1}{t(\log{t})^{5}}\right).$
@@
$\left|u_{k}(t)-\omega_{k}\right|$ $\lesssim\int_{t}^{\infty}\frac{ds}{s(\log{s})^{5}}$
$\lesssim\frac{1}{(\log{t})^{4}}.$
@@
$\begin{aligned} z_{k}(t)&=z_{0}(t)+\rho_{k}(t)u_{k}(t)\\ &=z_{\infty}+\rho_{k}(t)\omega_{k}+O\left(\frac{1}{(\log{t})^{3}}\right)\\ &=z_{\infty}+\omega_{k}\left(\log{t}-\frac{d-1}{2}\log{(\log{t})}+c_{\star}\right)+O\left(\frac{\log{(\log{t})}}{\log{t}}\right).\end{aligned}$ (6.35)
@@
Here, we prove Theorem 1.1.
@@
$\left\|u(t)-\left(Q(\cdot-z_{0}(t))-\sum_{k=1}^{3}Q(\cdot-z_{k}(t))\right)\right\|_{H^{1}}+\|{\partial}_{t}u(t)\|_{L^{2}}\lesssim t^{-1}.$
By choosing $c_{0}=c_{\star}$, the asymptotic behavior of $z_{0},z_{1},z_{2},z_{3}$ follows from Lemma 6.6. Hence Theorem 1.1 is proved.

### v2

@@ 6.1. Sharp estimates for lengths and angles
Since $L^{-3}$ is decreasing, (6.19) holds for all $t$ with $0<t-T_{1}\ll 1$, which contradicts the definition of $T_{1}$. Hence $T_{1}=\infty$. In this case, $\max{(d_{12},d_{13},d_{23})}$ is bounded and monotone increasing, so it converges to a nonzero limit; this contradicts (6.4). Therefore, we have for $t\gg 1$
$\max{(d_{12},d_{13},d_{23})}\lesssim\frac{1}{L^{3}}.$ (6.21)
By (6.6), (6.17) and (6.21), we obtain (6.5). This completes the proof.
@@ 6.2. Long-time dynamics of $(1,3)$-sign solitary waves
In the previous subsection, we refine the estimates for $\rho_{1},\rho_{2},\rho_{3}$. Using these improved bounds, we derive a more precise system of ODEs for $z_{0},z_{1},z_{2},z_{3}$.
@@
$\begin{aligned} \dot{z}_{0}&=-\frac{\mathcal{F}(L)}{L}(Z_{1}+Z_{2}+Z_{3})+O\left(\frac{1}{t(\log{t})^{3}}\right),\\ \dot{z}_{1}&=\frac{\mathcal{F}(L)}{L}Z_{1}+O\left(\frac{1}{t(\log{t})^{3}}\right),\\ \dot{z}_{2}&=\frac{\mathcal{F}(L)}{L}Z_{2}+O\left(\frac{1}{t(\log{t})^{3}}\right),\\ \dot{z}_{3}&=\frac{\mathcal{F}(L)}{L}Z_{3}+O\left(\frac{1}{t(\log{t})^{3}}\right).\end{aligned}$ (6.22)
@@
$\left|\mathcal{F}(\rho_{k})-\mathcal{F}(L)\right|\lesssim\frac{1}{t(\log{t})^{3}},\ \left|\frac{1}{\rho_{k}}-\frac{1}{L}\right|\lesssim\frac{1}{(\log{t})^{5}}.$
Substituting these bounds into the equations in Lemma 5.1 yields (6.22). ∎
@@
Let $\vec{u}$ be a (1,3)-sign solitary waves. Then, there exists $z_{g}\in\mathbb{R}^{d}$ such that we have for $t\gg 1$
$\left|\frac{z_{0}(t)+z_{1}(t)+z_{2}(t)+z_{3}(t)}{4}-z_{g}\right|\lesssim\frac{1}{(\log{t})^{2}}.$ (6.23)
@@
By Lemma 6.3, we have for $t\gg 1$
$\left|\frac{\dot{z}_{0}+\dot{z}_{1}+\dot{z}_{2}+\dot{z}_{3}}{4}\right|\lesssim\frac{1}{t(\log{t})^{3}}.$
Since the right-hand side is integrable, there exists $z_{g}\in\mathbb{R}^{d}$ such that $\frac{z_{0}+z_{1}+z_{2}+z_{3}}{4}$ converges to $z_{g}$ as $t\to\infty$. In particular, we have for $t\gg 1$
$\left|\frac{z_{0}(t)+z_{1}(t)+z_{2}(t)+z_{3}(t)}{4}-z_{g}\right|$ $\lesssim\int_{t}^{\infty}\frac{ds}{s(\log{s})^{3}}$
$\lesssim\frac{1}{(\log{t})^{2}}.$
@@
Next, we prove the convergence of $z_{0}$. The idea, as in [20, Lemma 4.8], is to argue that if $Z_{1}+Z_{2}+Z_{3}$ did not converge, then its magnitude would grow and eventually become much larger than $\rho_{12},\rho_{13},\rho_{23}$, leading to a contradiction.
@@
Let $\vec{u}$ be a (1,3)-sign solitary waves. Then, there exists $z_{\infty}\in\mathbb{R}^{d}$ such that we have for $t\gg 1$
$|z_{0}(t)-z_{\infty}|\lesssim\frac{1}{(\log{t})^{2}}.$ (6.24)
@@
$\begin{aligned} \dot{Z}_{1}&=\frac{\mathcal{F}(L)}{L}(2Z_{1}+Z_{2}+Z_{3})+O\left(\frac{1}{t(\log{t})^{3}}\right),\\ \dot{Z}_{2}&=\frac{\mathcal{F}(L)}{L}(Z_{1}+2Z_{2}+Z_{3})+O\left(\frac{1}{t(\log{t})^{3}}\right),\\ \dot{Z}_{3}&=\frac{\mathcal{F}(L)}{L}(Z_{1}+Z_{2}+2Z_{3})+O\left(\frac{1}{t(\log{t})^{3}}\right).\end{aligned}$ (6.25)
Moreover, we have for $t\gg 1$
$\dot{\rho}_{12}$ $=\frac{\mathcal{F}(L)}{L}\rho_{12}+O\left(\frac{1}{t(\log{t})^{3}}\right),$
$\dot{\rho}_{13}$ $=\frac{\mathcal{F}(L)}{L}\rho_{13}+O\left(\frac{1}{t(\log{t})^{3}}\right),$
$\dot{\rho}_{23}$ $=\frac{\mathcal{F}(L)}{L}\rho_{23}+O\left(\frac{1}{t(\log{t})^{3}}\right).$
Here, we define $\mathcal{R}$ as
$\mathcal{R}=\rho_{12}+\rho_{13}+\rho_{23}.$ (6.26)
Then, we have for $t\gg 1$
$\dot{\mathcal{R}}=\frac{\mathcal{F}(L)}{L}\mathcal{R}+O\left(\frac{1}{t(\log{t})^{3}}\right).$ (6.27)
Furthermore, we define $W=Z_{1}+Z_{2}+Z_{3}$ and $\mathcal{W}=|W|$. Then, $W$ satisfies
$\dot{W}=\frac{4\mathcal{F}(L)}{L}W+O\left(\frac{1}{t(\log{t})^{3}}\right).$
Therefore, there exists $C_{1}>1$ such that we have for $t\gg 1$
$\left|\dot{\mathcal{W}}-\frac{4\mathcal{F}(L)}{L}\mathcal{W}\right|\leq\frac{C_{1}\mathcal{F}(L)}{L^{3}}.$ (6.28)
@@
$\mathcal{W}(t^{\prime})>\frac{C_{1}}{L^{2}(t^{\prime})}.$ (6.29)
@@
Furthermore, we assume that $T_{1}<\infty$. Then, we have $\mathcal{W}(T_{1})=\frac{C_{1}}{L^{2}(T_{1})}$. Furthermore, by (6.28), we have
$\dot{\mathcal{W}}(T_{1})\geq\frac{4\mathcal{F}(L(T_{1}))}{L(T_{1})}\mathcal{W}(T_{1})-\frac{C_{1}\mathcal{F}(L(T_{1}))}{L^{3}(T_{1})}>0.$
Moreover, since $\frac{C_{1}}{L^{2}}$ is monotone decreasing, the estimate (6.30) also holds for $0<t-T_{1}\ll 1$, which contradicts the maximality of $T_{1}$. Hence $T_{1}=\infty$. Then, since $\mathcal{W}$ is monotone increasing for $t>t^{\prime}$, if $\mathcal{W}(t)$ did not tend to $\infty$ as $t\to\infty$, then $\mathcal{W}$ would converge. In that case, for $1\ll T_{2}$ we would have
$1\gtrsim\mathcal{W}(e^{T_{2}})-\mathcal{W}(T_{2})\gtrsim\int_{T_{2}}^{e^{T_{2}}}\frac{dt}{t\log t},$
@@
$=-\frac{3\mathcal{F}(L)}{L}\frac{\mathcal{R}}{\mathcal{W}}+O\left(\frac{\mathcal{W}+\mathcal{R}}{\mathcal{W}^{2}t(\log{t})^{3}}\right)$
$=-\frac{3\mathcal{F}(L)}{L}\mathcal{X}+O\left(\frac{1+\mathcal{X}}{t(\log{t})^{3}}\right),$
@@
which contradicts (6.33). Thus, we obtain
$\mathcal{W}=|z_{1}+z_{2}+z_{3}-3z_{0}|\lesssim\frac{1}{L^{2}}\sim\frac{1}{(\log{t})^{2}}.$
@@
We now prove Theorem 1.1. First, we establish the following lemma.
@@
Let $\vec{u}$ be a (1,3)-sign solitary waves. Then, there exist $z_{\infty}\in\mathbb{R}^{d}$ and equilateral triple $\omega_{1},\omega_{2},\omega_{3}\in S^{d-1}$ such that we have for $k=1,2,3$ and $t\gg 1$
@@
$\dot{Z}_{k}$ $=\mathcal{F}(\rho_{k})u_{k}+O\left(\frac{1}{t(\log{t})^{3}}\right),$
$\dot{\rho}_{k}$ $=\mathcal{F}(\rho_{k})+O\left(\frac{1}{t(\log{t})^{3}}\right).$
Moreover, we have
@@
$=O\left(\frac{1}{t(\log{t})^{4}}\right).$
@@
$\left|u_{k}(t)-\omega_{k}\right|$ $\lesssim\int_{t}^{\infty}\frac{ds}{s(\log{s})^{4}}$
$\lesssim\frac{1}{(\log{t})^{3}}.$
@@
$\begin{aligned} z_{k}(t)&=z_{0}(t)+\rho_{k}(t)u_{k}(t)\\ &=z_{\infty}+\rho_{k}(t)\omega_{k}+O\left(\frac{1}{(\log{t})^{2}}\right)\\ &=z_{\infty}+\omega_{k}\left(\log{t}-\frac{d-1}{2}\log{(\log{t})}+c_{\star}\right)+O\left(\frac{\log{(\log{t})}}{\log{t}}\right).\end{aligned}$ (6.35)
@@
Finally, we prove Theorem 1.1.
@@
$\left\|u(t)-\left(Q(\cdot-z_{0}(t))-\sum_{k=1}^{3}Q(\cdot-z_{k}(t))\right)\right\|_{H^{1}}+\|{\partial}_{t}u(t)\|_{L^{2}}\lesssim t^{-1},$
which is (1.6). Choosing $c_{0}=c_{\star}$, the asymptotic behaviors of $z_{0},z_{1},z_{2},z_{3}$ satisfy (1.7) by Lemma 6.6. This completes the proof of Theorem 1.1. ∎